# Low-Dimension-to-High-Dimension Generalization And Its Implications for Length Generalization

### v1

@@ Abstract
Low-Dimension-to-High-Dimension (LDHD) generalization is a special case of Out-of-Distribution (OOD) generalization, where the training data are restricted to a low-dimensional subspace of the high-dimensional testing space. Assuming that each instance is generated from a latent variable and the dimension of the latent variable reflects the problem scale, the inherent scaling challenge in length generalization can be captured by the LDHD generalization in the latent space. We theoretically demonstrate that LDHD generalization is generally unattainable without exploiting prior knowledge to provide appropriate inductive bias. Specifically, we explore LDHD generalization in Boolean functions. We verify that different architectures trained with (S)GD converge to min-degree interpolators w.r.t. different independent sets. LDHD generalization is achievable if and only if the target function coincides with this inductive bias. Applying the insights from LDHD generalization to length generalization, we explain the effectiveness of CoT as changing the structure latent space to enable better LDHD generalization. We also propose a principle for position embedding design to handle both the inherent LDHD generalization and the nuisances such as the data format. Following the principle, we propose a novel position embedding called RPE-Square that remedies the RPE for dealing with the data format nuisance.
@@ 1 Introduction
The field of learning to reason has gained significant popularity in the machine learning community due to its impressive performance in reasoning tasks such as natural language processing [16, 17], mathematics [8, 11], coding [28], symbolic logic [1, 9], and planning [30, 21]. Reasoning problems inherently require generalization beyond the training distribution, a challenge known as Out-of-Distribution (OOD) generalization [13, 27]. One of the most significant challenges in OOD generalization is scaling, also known as length generalization [3, 29], where models trained on small-scale instances must generalize to large-scale instances. Length generalization is crucial because the size of sample spaces often increases exponentially with the complexity of reasoning problems, leading to intractable sample complexity and computational costs for models that do not achieve length generalization.
Numerous works have investigated the length generalization problem in various tasks and proposed various techniques that help scaling in practice, including modifications to model architectures [20, 11, 12], transformations in data formats [14, 31], prompt engineering for Large Language Models (LLMs) Wei et al. [22], Feng et al. [7]. While some techniques work uniformly well across a wide class of problems, many are fragile and even ad-hoc, applicable only to specific problems with certain formats [32]. This can be largely attributed to the mismatch between the inherent problem scale and the formal scale in the input domain (e.g., the length of the input string in language models). We further elaborate the statement in Example 1.
@@
Consider the addition learning problem in the language domain, where the input is a string. For an $N$-digit-plus-$N$-digit (written as $N$-addition) addition of two numbers $x$ and $y$, where $x=x_{N-1}\dots x_{0}$, $y=y_{N-1}\dots y_{0}$, $x_{i},y_{i}\in\{0,\dots,9\}$ ($x_{N-1}>0$ or $y_{N-1}>0$), consider two formats of the instance: the Aligned and Reverse Format (ARF), where the instance is represented as “$\mathtt{x_{0}\dots x_{n-1}+y_{0}\dots y_{n-1}=}$”; and the Unaligned and Reverse Format (URF), where the instance is represented as “$\mathtt{x_{0}\dots x_{n_{x}-1}+y_{0}\dots y_{n_{y}-1}=}$”, $n_{c}=\operatorname*{arg\,max}_{i}\{c_{i}\neq 0\},c\in\{x,y\}$. The formal scale in the language domain is the length of the string. However, the length of the input strings in neither format faithfully reflects the inherent problem scale: in ARF, the length of the input strings is always invariant; in URF, a shorter string may correspond to a larger scale than a longer string (e.g., $s_{1}=$$1+1234=$” is $4$-addition, $s_{2}=$$123+123=$” is $3$-addition; however, $|s_{1}|<|s_{2}|$). Furthermore, we find in experiments that Relative Position Embedding (RPE) improves the length generalization performance of Transformers for ARF, while it provides little benefit for the length generalization of Transformers for URF.
To develop robust and transferable methods for scalable models, it requires a formulation that captures the inherent scaling challenge of length generalization and is invariant to the nuisances such as data format. To do so, we first notice that in the length generalization of a wide range of problems, the instances can be seen as generated from some latent space whose dimension reflects the scale.
@@
Let $\Sigma=\{0,\dots,9\}\times\{0,\dots,9\}$. The instance of $n$-addition $x_{n-1}\dots x_{0}+y_{n-1}\dots y_{0}$ can be represented by the state vector $h_{n}=\left[\left(x_{0},y_{0}\right),\dots,\left(x_{n-1},y_{n-1}\right)\right]$. The $k$-th dimension of $h_{n}$ corresponds to the $(k-1)$-th digits of the two addenda (i.e., $x_{k-1}$ and $y_{k-1}$). The expansion of the state vector in the dimension corresponds to the increase in the addendum digits. Therefore, the length generalization from $N_{0}$-addition to $N$-addition can be seen as a generalization from the low-dimensional latent space $\Sigma^{N_{0}}$ to the high-dimensional latent space $\Sigma^{N}$.
To catch the above intuition, we make the following assumption about the underlying data generation.
Assumption 1 (Data Generation in Length Generalization).
  1. 1.
    A latent variable $h$ is sampled from a subspace $\Sigma^{n}$ of dimension $n$;
  2. 2.
    The label $y$ is determined by the concept $c$ and the latent variable $h$, i.e., $y=c(h)$;
  3. 3.
    The hidden variable $h$ is transformed to an input sequence by the data format mapping $\phi$.
According to Assumption 1, the inherent scaling challenge can be captured by Low-Dimension-to-High-Dimension (LDHD) generalization, a special scenario of OOD generalization with structural distributional shifts. In an LDHD generalization problem, a model is trained with samples from a low-dimensional subspace while evaluated with samples from a higher-dimensional superspace. To further illustrate, we can consider that the training samples are from the space $\Sigma^{N_{0}}$ and the testing samples are from $\Sigma^{N}$, where $\Sigma$ is a field and the numbers $N_{0},N$ satisfy $N_{0}<N$. With slight abuse of notation, we also use $\Sigma^{n}$ to denote the embedding space of the subspace $\Sigma$ into a space $\Sigma^{\bar{N}}$ for some sufficiently large $\bar{N}\geq N>N_{0}$ (we use $\bar{N}=N$ below if there is no special statement).
@@
The main challenge of LDHD generalization is that the testing space has extra dimensions compared to the training space. As a result, the testing space contains instances with orthogonal components to the training space. The training samples can reveal no information about how these components contribute to the results. For instance, in Example 2, learning with $N_{0}$-addition solely does not tell how $(N_{0}+1)$-th digits to $N$-th digits contribute to the result unless we have the prior knowledge that the addition of each digit shares the same process. We formally this challenge as No-Free-Lunch Theorem of LDHD Generalization in Section 3.
@@
The No-Free-Lunch Theorem [25, 24] of LDHD Generalization necessitates the use of prior knowledge in the learning process in order to achieve LDHD generalization. In practice, the prior knowledge is usually incorporated via the inductive bias of the learning algorithms and the models. Abbe et al. [1] claimed that models with (S)GD converge to min-degree interpolators. However, the minimal degree-profile bias neither holds uniformly for all model architectures nor is sufficient for achieving LDHD generalization. To develop LDHD generalizable models, we investigate different architectures whose inductive bias with (S)GD can potentially go beyond the min-degree interpolators. While random feature models are shown to be min-degree interpolators, we prove that random feature models with projections, where inputs are transformed by projections before fed to the random feature models, converge to min-degree interpolators under different independent sets. We further consider Position-Only Linear Attentions with Advice (PLAA), i.e., linear attentions whose attention scores are determined only by positions, with additional hints about the scales of the instances. This model can be seen as a simplified abstraction of decoder-only transformers with a special focus on positional relations that are considered crucial for length generalization. We demonstrate the failure of PLAA models equipped with Absolute Position Embeddings (APEs) in general cases and show the PLAA models with Relative Position Embeddings (RPEs) are min-degree interpolators w.r.t. some independent sets. These results show that for models biased towards min-degree interpolators w.r.t. some independent sets, LDHD generalization can be achieved if and only if the target mapping is the min-degree interpolator w.r.t. the same set.
@@
The LDHD generalization perspective further provides insights on how Chain-of-Thought (CoT) [22] can help length generalization. We show that CoT can be seen as a change of the hidden space, where each dimension of the latent space is augmented with an additional middle state. This transformation could facilitate LDHD generalization in the latent space for some problems. Another implication is a principle of position embedding design for length generalization with Transformers: we need to handle both the inherent LDHD generalization and the nuisances such as the data format in the design of the position embeddings. Following this principle, we propose a novel position embedding named RPE-Square. The RPE-Square enhances the RPE with the ability to handle certain data format nuisances, evidently improving the length generalization of the URF addition over the Transformer with the RPE.
The main contributions of the paper are summarized as follows:
  • We propose to study LDHD generalization, which captures the inherent scaling challenge of a wide range of length generalization problems. LDHD generalization provides a formulation under which we can perform theoretical analysis for general length generalization problems.
  • We study the inductive bias of different model architectures under (S)GD, which can be exploited to incorporate prior knowledge to achieve LDHD generalization. In the context of Boolean functions, we show that different architectures under (S)GD can have inductive bias other than min-degree interpolators, which enables LDHD generalization for a wider class of problems.
@@
Notations. We use $[n]$ to represent the set of numbers $\{1,\dots,N\}$. We denote the set of all functions from the set ${\mathcal{X}}$ to the set ${\mathcal{Y}}$ as ${\mathcal{F}}_{{\mathcal{X}},{\mathcal{Y}}}$. We define $\operatorname{Proj}(x,V)$ as the coordinate of the projection of $x$ onto the space spanned by $V$ with the basis $V=\left[v_{1},\dots,v_{r}\right]$, i.e., $\left[\operatorname{Proj}(x,V)\right]_{i}=\langle x,v_{i}\rangle$ for all $i=1,\dots,r$. We use $A^{*}$ to denote the Kleene closure of the set $A$, i.e., $A^{*}=\bigcup_{k=0}^{\infty}A^{k}$. We use $\deg(p)$ to denote the degree of the polynomial $p$.
@@ 2 Related Work
Length generalization in reasoning problems. Length generalization is a key challenge in learning to reason, typically interpreted as the ability to learn with small-scale instances of a task and generalize to unseen large instances of the same task [3, 31]. Various reasoning tasks are considered to investigate length generalization, including arithmetic [11, 7], Boolean logic [1, 6], symbolic reasoning [29], etc. Despite the rich literature on length generalization on specific reason tasks, few works have considered challenges and overconditions of length generalization for general problems. The existing works that analyze general length generalization mainly focus on the change of the input sequence length [26, 2], which does not precisely capture the difficulties of length generalization. Our work proposes to consider LDHD generalization, which characterizes the inherent challenges of length generalization in many tasks.
Inductive Bias of Model Architectures and Algorithms. In the situation of learning with overparatermization or incomplete information, proper inductive bias is essential to select the true model from the hypothesis [15, 5]. One of the most studied scenarios is the inductive bias of different model architectures under (S)GD. Previous research shows (S)GD combined with different model architectures lead to different effects of implicit regularization, such as different norms of the learnable parameters Gunasekar et al. [10], Bartlett et al. [5] and different complexity characterizations of the models [18, 19, 1]. We specially mention Abbe et al. [1], which proposes that models trained with (S)GD are biased towards min-degree profile interpolators in the context of Boolean functions, which do not achieve general LDHD generalization. Our results show that the min-degree profile bias does not hold for all models. We further show that models with different architecture can converge to min-degree profile interpolators under different independent sets when trained with (S)GD. This partially explains how model architectures can affect length generalization in reasoning problems.
3 Low-Dimension-to-High-Dimension Generalization
@@
$\mathbb{E}_{X^{m}\sim{\mathcal{D}}_{1}^{m},X_{m+1}\sim{\mathcal{D}}_{2}}\left[\ell\left(\hat{f}_{X^{m},c}(X_{m+1}),c(X_{m+1})\right)\right]\leq\epsilon,$ (1)
@@
$\sum_{f\in{\mathcal{F}}/\left(c_{1}|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}({\mathcal{X}}_{N})}\left[\ell\left(c_{1}(x),f(x)\right)\right]=\sum_{f\in{\mathcal{F}}/\left(c_{2}|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}({\mathcal{X}}_{N})}\left[\ell\left(c_{2}(x),f(x)\right)\right].$
@@
Theorem 1 necessitates the consideration of structural assumptions on the concept class such that a learning algorithm could identify the target concept from the hypothesis with the imperfect information provided by the low-dimensional training data. For example, the concept class of linear classifiers with fixed weight vector on ${\mathcal{X}}={\mathbb{R}}^{d}$, i.e., ${\mathcal{C}}=\{\operatorname{sgn}\left(w_{0}^{\intercal}x+b\right)\mid w_{0}\in{\mathbb{R}}^{d},b\in{\mathbb{R}}\}$ with the $d^{\prime}$-dimensional training sample space ${\mathcal{X}}_{1}={\mathbb{R}}^{d^{\prime}}\times\{0\}^{d-d^{\prime}}$ and the $d$-dimensional testing sample space where $d^{\prime}<d$. For any concept $c\in{\mathcal{C}}$, a learning algorithm could not identify the true concept $c$ solely from the training data from ${\mathcal{X}}_{1}$ and the hypothesis of all linear classifiers. However, if a linear algorithm exploits the structure of the concept class that the weight vector always equals $w_{0}$, it can compute the target bias from the training data in ${\mathcal{X}}_{1}$ and thus identify the target concept $c$.
@@
We further investigate how model structures can affect LDHD generalization. We show in Section 4 that different models trained with (S)GD can be seen as min-degree interpolators under different functional bases in the context of Boolean functions, which is a joint inductive bias of the model structures and (S)GD. This suggests a principle for model design to achieve LDHD generalization: make the concept class “low-degree” under the independent set induced by the model structure.
@@
Many generalization challenges in reasoning problems can be attributed to LDHD generalization inherently. Length generalization is representative of these challenges and of significance in reasoning problems. Intuitively, length generalization means that a model trained with small-scale instances of a reasoning problem can perform well on large-scale instances of the problem. To formalize the intuition, we consider a typical instance being generated from a hidden state variable $h$ that represents the core of the instance and is transformed to the model input by a data format mapping. The hidden state space is of the form $\Sigma^{*}$ for some field $\Sigma$. The dimension $n$ of the subspace from which the hidden state variable is sampled, $n=\operatorname*{arg\,max}_{k}\{h\in\Sigma^{k},h\not\in\Sigma^{k+1}\}$ , reflects the increase in the scale of the problem. Besides, we define the concept class on the hidden space, depicting that the change of the data format does not change the intrinsic logic of the concepts. Assumption 1 describes the pipeline of the data generation in length generalization.

### v3

@@ Abstract
Low-Dimension-to-High-Dimension (LDHD) generalization, a subset of Out-of-Distribution (OOD) generalization, involves training on a low-dimensional subspace and testing in a high-dimensional space. Assuming instances are generated from latent variables reflecting problem scale, LDHD generalization captures the inherent scaling challenge of length generalization. We theoretically show that LDHD generalization is unattainable without appropriate inductive bias. Focusing on Boolean functions, we demonstrate that different architectures trained with (S)GD converge to min-degree interpolators w.r.t. different linearly independent sets, achieving LDHD generalization only when the target function aligns with this bias. From the perspective of LDHD generalization for length generalization, we explain the success of CoT in restructuring latent space for improved LDHD generalization. We further propose a principle for designing position embeddings to address both LDHD generalization and data format nuisances separately. Following the principle, we introduce RPE-Square, a novel embedding that enhances RPE to better handle data formats.
Machine Learning, ICML
@@ 1 Introduction
Learning to reason has gained significant popularity in the machine learning community due to its impressive performance in reasoning tasks such as natural language processing (OpenAI, 2023a, b), mathematics (Frieder et al., 2023; Jelassi et al., 2023), coding (Zhang et al., 2022a), symbolic logic (Abbe et al., 2023; Garcez et al., 2022), and planning (Zhao et al., 2023; Valmeekam et al., 2023). One of the most significant challenges in learning to reason is length generalization (Anil et al., 2022; Zhang et al., 2022b), where models trained on small-scale instances must generalize to large-scale instances. Length generalization is crucial because the size of sample spaces often increases exponentially with the complexity of reasoning problems, leading to intractable sample complexity and computational costs for models that do not achieve length generalization.
Numerous works have investigated various techniques for length generalization, including modifications to model architectures (Shaw et al., 2018; Jelassi et al., 2023; Kazemnejad et al., 2024), transformations in data formats (Lee et al., 2023; Zhou et al., 2023), prompt engineering for Large Language Models (LLMs) (Wei et al., 2022; Feng et al., 2024). Although some of the above techniques work uniformly well across a wide class of problems, many are fragile and even ad-hoc, applicable only to specific problems with certain formats (Zhou et al., 2024). This is due to the mismatch between the inherent problem scale and the length of the input string. We illustrate it in the next Example 1.
@@
Consider the addition learning problem where the input is a string. For an NN-digit-plus-NN-digit (written as $N$-addition) addition of two numbers $x$ and $y$, where $x=x_{N-1}\dots x_{0}$, $y=y_{N-1}\dots y_{0}$, $x_{i},y_{i}\in\{0,\dots,9\}$ ($x_{N-1}>0$ or $y_{N-1}>0$), consider two formats: the Aligned and Reverse Format (ARF), where the instance is represented as $\mathtt{x_{0}\dots x_{n-1}+y_{0}\dots y_{n-1}=}$”; and the Unaligned and Reverse Format (URF), where the instance is represented as $\mathtt{x_{0}\dots x_{n_{x}-1}+y_{0}\dots y_{n_{y}-1}=}$”, $n_{c}=\operatorname*{arg\,max}_{i}\{c_{i}\neq 0\},c\in\{x,y\}$. However, the length of the input strings in neither format faithfully reflects the problem scale: in ARF, the length of the input strings is always invariant; in URF, a shorter string may correspond to a larger scale than a longer string (e.g., $s_{1}=$1+1234=1+1234=” is $4$-addition, $s_{2}=$123+123=123+123=” is $3$-addition; $s_{1}$ is of larger scale than $s_{2}$ but $\operatorname{length}(s_{1})<\operatorname{length}(s_{2})$).
(a) In-Distribution Generalization.
(b) (Typical) OOD Generalization.
(c) LDHD Generalization.
Figure 1: Illustrative comparison of in-distribution generalization, typical OOD generalization, and LDHD generalization. (a) In-distribution generalization assumes identical training and testing distributions. (b) Typical OOD generalization involves a shift between training and testing distributions, which remain relatively “close” (e.g., sharing support or having small distributional distances). (c) LDHD generalization features a training distribution restricted to a low-dimensional subspace and a testing distribution on a high-dimensional space, often vastly different. While LDHD is a type of OOD generalization, its structured shift poses unique challenges, as training data provide no clues about the additional dimensions’ contribution to the label.
Example 1 demonstrates the sensitivity of length generalization to the data format. Developing robust and transferable methods for scalable models requires a formulation that captures the inherent scaling challenge of length generalization and is invariant to the data format nuisance.
1.1 Our Main Works
Low-Dimension-to-High-Dimension Generalization Perspective for Length Generalization. We propose 1 that disentangles the problem scale and the data format, which provides a more precise formulation for the analysis of length generalization.
Abstraction 1 (Data Generation in Length Generalization).
1. A latent variable $h$ is sampled from a subspace $\Sigma^{n}$ of dimension $n$;
2. The label $y$ is determined by the concept $c$ and the latent variable $h$, i.e., $y=c(h)$;
3. The latent variable $h$ is transformed to an input sequence by the data format mapping $\phi$.
In 1, the scale of the instance is captured by the dimension of the latent space from which the instance is sampled. The scale shift in length generalization is characterized by Low-Dimension-to-High-Dimension (LDHD) generalization in the latent space. In other words, the generalization from “short” instances to “long” instances corresponds to the generalization from the low-dimension latent subspace of the training data to the high-dimension latent subspace. The computation of the task is represented by the concept $c$, which is defined on the latent space and independent of the data format. The data format decides how a hidden variable is mapped to an actual input sequence. The next Example 2 illustrates the abstraction in the addition.
@@ 1. A latent variable $h$ is sampled from a subspace $\Sigma^{n}$ of dimension $n$;
Let $\Sigma=\{0,\dots,9\}^{2}$. The instance of $n$-addition $x_{n-1}\dots x_{0}+y_{n-1}\dots y_{0}$ can be represented by the latent vector $h_{n}=\left[\left(x_{0},y_{0}\right),\dots,\left(x_{n-1},y_{n-1}\right)\right]$. The $k$-th dimension of $h_{n}$ corresponds to the (k1)(k-1)-th digits of the two addenda (i.e., $x_{k-1}$ and $y_{k-1}$). The expansion of the state vector in the dimension corresponds to the increase in the addendum digits. In the length generalization of the addition task, the scale shift from $N_{0}$-addition to $N$-addition can be seen as a generalization from the low-dimensional latent space $\Sigma^{N_{0}}$ to the high-dimensional latent space $\Sigma^{N}$. The label, i.e., the sum of the addenda, is determined by the computation of the addition task. The input string is determined by both the latent variable and the data format (e.g., ARF or URF).
Example 2 demonstrates that each instance of $n$-addition is generated from an $n$-dimension latent variable via a data formant mapping. Length generalization of addition requires addressing both LDHD generalization in the latent space, which captures the computation of addition as the scale increases, and the data format, which may hinder the model in learning the addition operator. In general, from the LDHD generalization perspective, length generalization involves two key aspects: LDHD generalization in the latent space and the nuisance of the data format. LDHD generalization reflects the scale shift between the training and test data, which is the inherent challenge of length generalization.
@@
No-Free-Lunch Theorem of LDHD Generalization. The main challenge of LDHD generalization is that the testing space has extra dimensions compared to the training space. As a result, the testing space contains instances with orthogonal components to the training space. The training samples cannot reveal any information about how these components contribute to the results. For instance, in Example 2, learning with $N_{0}$-addition solely does not tell how $(N_{0}+1)$-th digits to $N$-th digits contribute to the result unless we have the prior knowledge that the addition of each digit shares the same process. Figure 1 shows an intuitive illustration for the challenge of LDHD generalization of linear models on ${\mathbb{R}}^{2}$, compared to in-distribution generalization and typical OOD generalization settings. When learning a linear decision boundary from the low-dimension training data, we need to know the slope of the decision boundary as a prior; otherwise, we will fail to achieve LDHD generalization: While all the three solid lines in Figure 1(c) perfectly separate the low-dimension training data, only the true decision boundary can achieve LDHD generalization.
We formalize the above challenge as No-Free-Lunch Theorem (Wolpert & Macready, 1997; Wolpert, 2002) of LDHD Generalization. Theorem 1 shows that no algorithm could achieve LDHD generalization uniformly for all tasks, which necessitates the use of prior knowledge in the learning process in order to achieve LDHD generalization.
@@
Inductive Bias of Models and Algorithms for LDHD Generalization. In practice, the prior knowledge is usually incorporated via the inductive bias of the learning algorithms and the models. To develop LDHD generalizable models, we investigate the inductive bias of different architectures trained by (S)GD. While random feature models are shown to be min-degree interpolators (Abbe et al., 2023), Theorem 2 shows that random feature models with projections, where inputs are transformed by projections before fed to the random feature models, converge to min-degree interpolators w.r.t. different linearly independent sets. Furthermore, we consider Position-Only Linear Attentions with Advice (PLAA), i.e., linear attentions whose attention scores are determined only by positions, with additional hints about the scales of the instances. This model can be seen as a simplified abstraction of decoder-only transformers with a special focus on positional relations that are considered crucial for length generalization. Theorems Theorem 3 and Theorem 4 show that PLAA with Absolute Position Embedding (APE) and PLAA with Relative Position Embedding (RPE) converge to min-degree interpolators w.r.t. different linearly independent sets. The results illustrate the limitation of PLAA with APE in LDHD generalization and the potential of PLAA with RPE to overcome the limitation.
@@
Implications of the LDHD Generalization Perspective for Practical Length Generalization Techniques. The LDHD generalization perspective further provides insights into practical length generalization techniques. In Section 5.1, we discuss the role of Chain-of-Thought (CoT) (Wei et al., 2022) in length generalization. We show that CoT can be seen as a change of the hidden space, where each dimension of the latent space is augmented with an additional middle state. This transformation could facilitate LDHD generalization in the latent space for some problems. Additionally, in Section 5.2, we propose a principle of position embedding design for length generalization with Transformers: We need to handle both the inherent LDHD generalization and the nuisances such as the data format in the design of the position embeddings. Following this principle, we propose a novel position embedding named RPE-Square. Our experiments show that the RPE-Square evidently enhances the RPE with the ability to handle the nuisance of the unaligned data format.
The next sections of the paper are organized as follows: In Section 3, we introduce the formal definition and No-Free-Lunch Theorem of LDHD generalization; in Section 4, we present the results of the inductive bias of different models trained by gradient and how the inductive bias influences LDHD generalization; in Section 5, we show how the LDHD perspective helps to understand and design practical techniques for length generalization.
1.2 Notations
@@
We use $[n]$ to represent the set of numbers $\{1,\dots,N\}$. We denote the set of all functions from the set $\mathcal{X}$} to the set $\mathcal{Y}$} as ${\mathcal{F}}_{{\mathcal{X}},{\mathcal{Y}}}$. We define $\operatorname{Proj}(x,V)$ as the coordinate of the projection of $x$ onto the space spanned by $V$ with the basis $V=\left[v_{1},\dots,v_{r}\right]$, i.e., $\left[\operatorname{Proj}(x,V)\right]_{i}=\langle x,v_{i}\rangle$ for all $i=1,\dots,r$. We use $A^{*}$ to denote the Kleene closure of the set $A$, i.e., $A^{*}=\bigcup_{k=0}^{\infty}A^{k}$. We use $\deg(p)$ to denote the degree of the polynomial $p$. We represent the set of $N\times N$ upper triangular matrix as ${\mathcal{U}}_{N}$.
@@ 2 Related Work
Length generalization in reasoning problems. Length generalization is a key challenge in learning to reason, typically interpreted as the ability to learn with small-scale instances of a task and generalize to unseen large instances of the same task (Anil et al., 2022; Zhou et al., 2023). Various reasoning tasks are considered to investigate length generalization, including arithmetic (Jelassi et al., 2023; Feng et al., 2024), Boolean logic (Abbe et al., 2023; d’Ascoli et al., 2023), symbolic reasoning (Zhang et al., 2022b), etc. Despite the rich literature on length generalization on specific reason tasks, few works have considered challenges and overconditions of length generalization for general problems. The existing works that analyze general length generalization mainly focus on the change of the input sequence length (Xiao & Liu, 2023; Ahuja & Mansouri, 2024; Huang et al., 2024). This formulation conflates the inherent scaling challenges with the effects of data format, making it difficult to precisely capture the challenges of length generalization. To address this, our work proposes disentangling length generalization into two distinct aspects: LDHD generalization in the latent space, which characterizes the challenges associated with increasing scale, and the data format nuisance.
Inductive Bias of Model Architectures and Algorithms. In the situation of learning with overparatermization or incomplete information, proper inductive bias is essential to select the true model from the hypothesis (Neyshabur, 2017; Bartlett et al., 2021; Teney et al., 2024). One of the most studied scenarios is the inductive bias of different model architectures under (S)GD. Previous research shows (S)GD combined with different model architectures lead to different effects of implicit regularization, such as different norms of the learnable parameters Gunasekar et al. (2017); Bartlett et al. (2021) and different complexity characterizations of the models (Razin & Cohen, 2020; Razin et al., 2021; Abbe et al., 2023). We specially mention Abbe et al. (2023), which proposes that models trained with (S)GD are biased towards min-degree profile interpolators in the context of Boolean functions, which do not achieve general LDHD generalization. Our results show that the min-degree profile bias does not hold for all models. We further show that models with different architecture can converge to min-degree profile interpolators under different linearly independent sets when trained with (S)GD. This partially explains how model architectures can affect length generalization in reasoning.
3 LDHD Generalization
@@
$\mathbb{E}_{X^{m}\sim{\mathcal{D}}_{1}^{m},X_{m+1}\sim{\mathcal{D}}_{2}}\left[\ell\left(\hat{f}_{X^{m},c}(X_{m+1}),c(X_{m+1})\right)\right]\leq\epsilon,$
@@
$=$ $\sum_{f\in{\mathcal{F}}/\left(c_{2}|_{{\mathcal{X}}_{N_{0}}}\right)}\mathbb{E}_{x\sim{\mathcal{D}}({\mathcal{X}}_{N})}\left[\ell\left(c_{2}(x),f(x)\right)\right].$
@@
Theorem 1 necessitates the consideration of structural assumptions on the concept class such that a learning algorithm could identify the target concept from the hypothesis with the imperfect information provided by the low-dimensional training data. For example, the concept class of linear classifiers with fixed weight vector on ${\mathcal{X}}={\mathbb{R}}^{d}$, i.e., ${\mathcal{C}}=\{\operatorname{sgn}\left(w_{0}^{\intercal}x+b\right)\mid w_{0}\in{\mathbb{R}}^{d},b\in{\mathbb{R}}\}$ with the $d^{\prime}$-dimensional training sample space ${\mathcal{X}}_{1}={\mathbb{R}}^{d^{\prime}}\times\{0\}^{d-d^{\prime}}$ and the $d$-dimensional testing sample space where $d^{\prime}<d$. For any concept $c\in{\mathcal{C}}$, a learning algorithm could not identify the true concept $c$ solely from the training data from ${\mathcal{X}}_{1}$ and the hypothesis of all linear classifiers. However, if a learning algorithm exploits the structure of the concept class, i.e., by fixing the weight vector to $w_{0}$ in prior, it can compute the target bias from the training data in ${\mathcal{X}}_{1}$ and then identify the target concept $c$.
@@
We further investigate how model structures can affect LDHD generalization. We show in Section 4 that different models trained with (S)GD can be seen as min-degree interpolators under different functional bases in the context of Boolean functions, which is a joint inductive bias of the model structures and (S)GD. This insight suggests a principle for model design to achieve LDHD generalization: ensure the concept class is “low-degree” under the linearly independent set induced by the model structure.
@@
LDHD generalization captures the challenge of scale shift in length generalization. Intuitively, length generalization means that a model trained with small-scale instances of a reasoning problem can perform well on large-scale instances of the problem. To formalize the intuition, we consider a typical instance being generated from a hidden variable $h$ that represents the “core” of the instance and is transformed to the model input by a data format mapping. The hidden space is of the form $\Sigma^{*}$ for some domain $\Sigma$. The dimension $n$ of the subspace from which the hidden variable is sampled, $n=\operatorname*{arg\,max}_{k}\{h\in\Sigma^{k},h\not\in\Sigma^{k+1}\}$ , reflects the increase in the scale of the problem. Besides, we define the concept class on the hidden space, depicting that the change of the data format does not change the computation of the concepts. 1 describes the pipeline of the data generation in length generalization.
Remark 1.
Some works observe length generalization on certain tasks without domain-specific priors, using only a small fraction of large-scale data during pretraining. This does not contradict our No-Free-Lunch Theorem, as “no” and “few” large-scale examples are fundamentally different. Prior work (Jelassi et al., 2023) shows that exposure to a tiny fraction of long sequences in pre-training, which is called priming, can significantly improve length generalization. From our perspective, this is because the presence of long sequences, even in small quantities, prevents the model from learning an overly simple interpolator that fails to extrapolate. For example, in the addition task, training only on 3-digit numbers may lead the model to ignore digits beyond the third, while including just a few 5-digit examples compels it to process later digits, enhancing extrapolation.